OpenStudy (anonymous):

Factor? (2A+3B)(2A-3B)

OpenStudy (saifoo.khan):

isnt it already factored?

OpenStudy (anonymous):

thats already factored form .

OpenStudy (saifoo.khan):

Yeah!

OpenStudy (anonymous):

No, multiply.....to get 4a^2 - 9b^2

OpenStudy (anonymous):

identity (a - b) * (a + b) = a² - b² So, (2a+3b)(2a-3b) = (2a)² - (3b)² = 4a² - 9b² => 4a to the second power - 9b to the second power